OpenStudy (anonymous):

solve by the elimination method x+7y=58 -x+3y=22

OpenStudy (anonymous):

\[3y-22 = x\] \[3y-22+7y=58\] \[10y = 80\] \[y = 8\] \[x + (8*7) = 58\] \[x= 58 - 56\] \[x = 2\]